OpenStudy (anonymous):

Solving Cubic Equations By Factoring: X^3+27=0

OpenStudy (anonymous):

i think its (x+3)(x^2-3x+9)=0

OpenStudy (anonymous):

yea thats it, so now set x+3=0 and (x^2-3x+9)=0 and solve for x in both

OpenStudy (anonymous):

-3 because x^3=-27 Take cube root of both side x= cube root of (-27)=-3

OpenStudy (anonymous):

Brie, you would lose some of the solutions by doing that. this problem has more than 1 x

OpenStudy (anonymous):

I have this: x+3=0 and (x^2-3x+9)=0 but how do you do it after this point?

OpenStudy (anonymous):

so for the left one you just subtract 3 and get x=-3....for the right one you would either use the quadratic equation or factor it first

OpenStudy (anonymous):

Okay, Thank You!

OpenStudy (anonymous):

np, but you would get imaginary numbers for the second one and I dont know if your teacher is looking for that or not